Project

General

Profile

01022014 Minutes

Attending: Mat, Minnesota

Outfitting:
Fixing diblock 4 cooling issues.
12/31-1/1 - The low hit rate is more heavily populated on onmon spectra. (APD work on the detector)

Production:
No new data processed: Since before 12/15, 12920
-No raw files transfered to disk.
Files now going to enstore.